OpenStudy (anonymous):

Find the function h(x) = f(x) - g(x) if f(x) = 3x and g(x) = 32x - 3x. h(x) = 0 h(x) = -32x h(x) = 2(32x) h(x) = 3x(2 - 3x)

OpenStudy (texaschic101):

h(x) = f(x) - g(x) h(x) = 3x - 32x - 3x h(x) = -32x --- the 3x's cancel out

OpenStudy (anonymous):

what is the answer

OpenStudy (cathyangs):

@texaschic101 showed you the answer! Their last line: "h(x) = -32x --- the 3x's cancel out". the answer is h(x) = -32x